# Robust First and Second-Order Differentiation for Regularized Optimal Transport

### v1

@@ 4.2 Coupling matrices with uniform marginal distributions
The largest singular value of $\Pi$ is $\sigma_{1}=\frac{1}{\sqrt{NM}}$ and it is simple. Hence, by Prop.8, the largest and smallest eigenvalues of $H(\Pi)$ are $\lambda_{1}(H)=\frac{1}{M}+\frac{1}{N}$, $\lambda_{M+N}(H)=0$, and both are simple.
The second smallest eigenvalue of $H$, denoted by $\lambda_{M+N-1}(H)$, can be obtained from $\left(\lambda-\frac{1}{M}\right)\left(\lambda-\frac{1}{N}\right)=\sigma_{2}^{2}$, i.e., $\lambda^{2}-\lambda_{1}\lambda+\sigma_{1}^{2}-\sigma_{2}^{2}=0$. With $\Delta:=\frac{(\sigma^{2}_{1}-\sigma^{2}_{2})}{\lambda_{1}^{2}}$, this gives
$\lambda_{M+N-1}(H)$ $=\frac{1}{2}\lambda_{1}\left[1-\sqrt{1-4\Delta}\right]=\frac{1}{2}\lambda_{1}\frac{4\Delta}{1+\sqrt{1-4\Delta}}.$ (52)
Hence, using the fact that $1\leq 1+\sqrt{1-4\Delta}\leq 2$ we obtain $\frac{(\sigma^{2}_{1}-\sigma^{2}_{2})}{\lambda_{1}}\leq\lambda_{M+N-1}(H)\leq\frac{2(\sigma^{2}_{1}-\sigma^{2}_{2})}{\lambda_{1}}$. To obtain the bounds for the condition numbers, by (52), we have
$\frac{(M+N)^{2}}{2M^{2}N^{2}(\sigma_{1}^{2}-\sigma_{2}^{2})}=\frac{\lambda_{1}^{2}}{2(\sigma^{2}_{1}-\sigma^{2}_{2})}\leq\kappa(H)\leq\frac{\lambda_{1}^{2}}{(\sigma^{2}_{1}-\sigma^{2}_{2})}=\frac{(M+N)^{2}}{M^{2}N^{2}(\sigma_{1}^{2}-\sigma_{2}^{2})}.$
This gives the bounds in (51). ∎
@@
Let $M=N$ and $\Pi$ be a positive coupling matrix with uniform marginals. The eigenvalues of $H(\Pi)$ are
$\lambda_{j}(H)=\frac{1}{N}+\sigma_{j},\quad\lambda_{2N+1-j}(H)=\frac{1}{N}-\sigma_{j},\quad 1\leq j\leq N,$ (53)
where $\{\sigma_{j}\}$ are the singular values of $\Pi$ in descending order. In particular, $\sigma_{1}=\frac{1}{N}$, $\lambda_{1}(H)=\frac{2}{N}$ and $\lambda_{2N}=0$. The condition number of $H(\Pi)$ is bounded by
$\frac{2}{N^{2}(\sigma_{1}^{2}-\sigma_{2}^{2})}\leq\kappa(H)\leq\frac{4}{N^{2}(\sigma_{1}^{2}-\sigma_{2}^{2})}.$ (54)
4.3 Condition number of H-matrices in entropy-regularized OT
We establish in this section lower and upper bounds for the condition number of the H-matrix when $\Pi^{*}$ is approximated by the Sinkhorn algorithm for uniform marginal distributions.
The Sinkhorn algorithm alternatively re-scales the rows and columns of the coupling matrix to achieve the marginal constraints. They produce a sequence of coupling matrices $\{\Pi^{(l)}\}$ that converges to $\Pi^{*}$ entry-wisely, i.e., $\lim_{l\rightarrow+\infty}\Pi_{ij}^{(l)}=\Pi_{ij}^{*}$ [29, 30, 21]. In practice, the Sinkhorn iteration stops when a criterion is met. One stopping criterion is that the marginal distributions of $\Pi^{(l)}$ are entry-wise $\delta$ away from the given $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$. Thus, an important question is whether the condition number of $H(\Pi^{(l)})$ is controlled.
The next proposition shows that if $\delta=\|\Pi-\Pi^{*}\|_{F}$ is small with $\|\cdot\|_{F}$ denotes the Frobenius norm, the condition number of the H-matrix of $\Pi^{(l)}$ is almost as large as the condition number of $H(\Pi^{*})$. The proof is based on Weyl’s inequality, and we postpone it to Appendix A.
Proposition 11 (Condition number of H-matrix in Sinkhorn).
Let $\Pi^{*}$, with uniform marginal distributions, be the optimal coupling matrix minimizing the EOT distance. Assume that the coupling matrix $\widehat{\Pi}$ is computed by an early-stopped Sinkhorn algorithm that satisfies
@@ 4.3 Condition number of H-matrices in entropy-regularized OT
Then, the eigenvalues of $H(\widehat{\Pi})$ satisfies
$|\lambda_{k}(H(\widehat{\Pi}))-\lambda_{k}(H(\Pi^{*}))|\leq\delta+\delta_{2},\quad 1\leq k\leq N+M.$ (56)
In particular, if $\delta+\delta_{2}=t\frac{MN}{M+N}(\sigma_{1}^{2}-\sigma_{2}^{2})$ with $t\in[0,1)$, where $\sigma_{1},\sigma_{2}$ denoting the largest two singular values of $\Pi^{*}$, the condition number of $H(\widehat{\Pi})$ is bounded by
$\frac{1-t\Delta}{(2+t)\Delta}\leq\kappa(H(\widehat{\Pi}))\leq\frac{1+t\Delta}{(1-t)\Delta},$
where $\Delta=(\frac{MN}{M+N})^{2}(\sigma_{1}^{2}-\sigma_{2}^{2})$, while $\frac{1}{2\Delta}\leq\kappa(H(\Pi^{*}))\leq\frac{1}{\Delta}$.
The above bounds apply to general H-matrices in entropy-regularized Sinkhorn algorithms. However, these bounds do not show explicit dependence on $\epsilon$, the strength of regularization. In the next section shows, we study the dependence of the condition number on $\epsilon$ and $N$ for point-clouds datasets.
4.4 Ill-conditioned H-matrices from data clouds
We investigate in this section the condition number of the $H$-matrix for a specific example of EOT that matches data clouds with $N$ points. In this simple setting, $M=N$ and both marginals are uniform, so the condition number of the $H$-matrix is $\kappa=\frac{2}{N\lambda_{2N-1}}$ by By Corollary 10. Therefore, it suffices to investigate the dependence of smallest positive eigenvalue, $\lambda_{2N-1}$, on $N$ and $\epsilon$.
We show that the smallest positive eigenvalue of the $H$-matrix can decay at rate $O(e^{-\frac{1}{\epsilon}})$ for a fixed $N$ and at $O(1/N)$ for a fixed $\epsilon$. These asymptotic results are proved for equally-spaced points on the unit circle in Example 12, and are numerically demonstrated for random data clouds sampled from a uniform distribution.
@@ 4.4 Ill-conditioned H-matrices from data clouds
Consider $N$ equally spaced points on the unit circle $\{\boldsymbol{y}_{i}=\left[\begin{array}[]{ccccccccc}\cos x_{i}&\sin x_{i}\end{array}\right]\}_{i=0}^{N-1}$, where $x_{i}=\frac{2\pi i}{N}$. Let $\boldsymbol{\mu}=\frac{1}{N}\mathbbm{1}_{N}$ be the uniform distribution, we are interested in the spectrum of $H$ associate to the symmetric entropic regularized optimal transport loss $\text{OT}_{\epsilon}(C^{\boldsymbol{Y}\rightarrow\boldsymbol{Y}},\boldsymbol{\mu},\boldsymbol{\mu})$. The coupling matrix is
$\Pi^{*}=\underset{\Pi\in\mathbb{R}^{N\times N}_{\geq 0}:\Pi\mathbbm{1}_{N}={\boldsymbol{\mu}},\Pi^{\top}\mathbbm{1}_{N}={\boldsymbol{\mu}}}{\operatorname{arg}\operatorname{min}}\;\sum_{i=1}^{N}\sum_{j=1}^{N}C_{ij}\Pi_{ij}+\epsilon\text{KL}(\Pi,{\boldsymbol{\mu}}\otimes\boldsymbol{\mu})$
with $\epsilon>0$, where $C_{ij}=\|\boldsymbol{y}_{i}-\boldsymbol{y}_{j}\|_{2}^{2}$. Also, let $H:=H(\Pi^{*})$ and denote its condition number by $\kappa(H)=\frac{\lambda_{1}(H)}{\lambda_{2N-1}(H)}$. Then, the following statements hold true.
@@
    $\Pi^{*}=\frac{K}{\lambda_{1}(K)N}$, where the Gibbs kernel $K\in\mathbb{R}^{N\times N}$ is a symmetric matrix with entries $K_{ij}=\exp\left(-\frac{C_{ij}}{\epsilon}\right)$ and $\lambda_{1}(K)$ is the largest eigenvalue of $K$.
@@
    The first two singular values of $\Pi^{*}$ are $\sigma_{1}(\Pi^{*})=\frac{1}{N}$, $\sigma_{2}(\Pi^{*})=\frac{\lambda_{2}(K)}{\lambda_{1}(K)N}$, and the largest and smallest positive eigenvalues of $H$ are $\lambda_{1}(H)=\frac{2}{N}$, and $\lambda_{2N-1}(H)=\frac{1}{N}-\sigma_{2}(\Pi^{*})$, where $\lambda_{2}(K)$ is the second largest eigenvalue of $K$.
@@
    The smallest positive eigenvalue of $H$ and the condition number of $H$ satisfies
    $\lim_{\epsilon\rightarrow 0^{+}}\lim_{N\rightarrow+\infty}\frac{N\cdot\lambda_{2N-1}(H)}{\epsilon}$ $=\frac{1}{4},\quad$ $\lim_{\epsilon\rightarrow 0^{+}}\lim_{N\rightarrow+\infty}\epsilon\cdot\kappa(H)$ $=8,$ (57)
    $\lim_{N\rightarrow+\infty}\lim_{\epsilon\rightarrow 0^{+}}\frac{N\lambda_{2N-1}(H)}{r_{N,\epsilon}}$ $=4\pi^{2},\quad$ $\lim_{N\rightarrow+\infty}\lim_{\epsilon\rightarrow 0^{+}}r_{N,\epsilon}\kappa(H)$ $=\frac{1}{2\pi^{2}},$ (58)
@@
We postpone the proof to Appendix A.
For a fixed $\epsilon$, when $N$ is large enough, the smallest positive eigenvalue $\lambda_{2N-1}$ scales as $\frac{\epsilon}{4N}$, which is numerically illustrated in Figure 1(a). Then the condition number scales as $\frac{8}{\epsilon}$. Meanwhile for fixed $N$, when $\epsilon$ is small enough (e.g., when $\epsilon<\frac{4\pi^{2}}{N^{2}}$), the smallest positive eigenvalue scales as $\frac{4\pi^{2}r_{N,\epsilon}}{N}$, which is numerically illustrated in Figure 1(b). Then the condition number scales as $\frac{1}{2\pi^{2}r_{N,\epsilon}}$, which grows exponentially at rate $O(e^{-\frac{1}{\epsilon}})$. For example, for $N=50$ and $\epsilon=0.0001$, the condition number of $H$ is larger than $10^{70}$; in this case, a truncated SVD for $H$ is crucial in calculating the Hessian of EOT and the gradient of Sinkhorn distance. On the other hand, when $N=1500$ and $\epsilon=0.0001$, the condition number of $H$ is only about $8\times 10^{4}$. In addition, this scale phenomenon is also observed for some random datasets as well.
Figure 1: Decay of the smallest positive eigenvalue $\lambda_{2N-1}$ in $N$ and $\epsilon$. Equally spaced points on the unique circle: (a) $\lambda_{2N-1}\approx\frac{\epsilon}{4N}$ when $N>\frac{2\pi}{\sqrt{\epsilon}}$; (b) $\lambda_{2N-1}\approx 4\pi^{2}r_{N,\epsilon}$ when $\epsilon<\frac{4\pi^{2}}{N^{2}}$. Uniformly distributed point cloud in unit square: (c) $\lambda_{2N-1}=O(\frac{1}{N})$ when $N$ is large; (d) $\lambda_{2N-1}=O(e^{-\frac{1}{\epsilon}})$ when $\epsilon$ is small.
Next, we further numerically investigate the case of point-clouds datasets that are sampled from the uniform distribution in the unit square $[0,1]^{2}$. Similarly, we are interested in the spectrum of $H$ associated with $\text{OT}_{\epsilon}(C^{\boldsymbol{Y}\rightarrow\boldsymbol{Y}},\boldsymbol{\mu},\boldsymbol{\mu})$. In figure 1(d) show that $\lambda_{2N-1}=O(e^{-\frac{1}{\epsilon}})$ when $\epsilon$ is small enough for each fixed N, and Figure 1(c) show $\lambda_{2N-1}=O(\frac{1}{N})$ when $N$ is large for each fixed $\epsilon$. These asymptotic orders are the same as the analytical results proved in Example 12, but the exact limit depends on the distribution of data points, and it is beyond the scope of this study.
In summary, the $H$-matrix can be severely ill-conditioned with the smallest eigenvalue at the order of $O(e^{-\frac{1}{\epsilon}})$ when $\epsilon$ is small, or $O(\frac{1}{N})$ when $N$ is large. Thus, when solving a linear system with $H$, it is important to properly regularize the ill-posed inverse problem.

### v2

@@ 4.2 Coupling matrices with uniform marginal distributions
The largest singular value of $\bm{\Pi}$ is $\sigma_{1}=\frac{1}{\sqrt{NM}}$ and it is simple. Hence, by Prop.9, the largest and smallest eigenvalues of $\mathbf{H}$ are $\lambda_{1}(\mathbf{H})=\frac{1}{M}+\frac{1}{N}$, $\lambda_{M+N}(\mathbf{H})=0$, and both are simple.
The second smallest eigenvalue of $\mathbf{H}$, denoted by $\lambda_{M+N-1}(\mathbf{H})$, can be obtained from $\left(\lambda-\frac{1}{M}\right)\left(\lambda-\frac{1}{N}\right)=\sigma_{2}^{2}$, i.e., $\lambda^{2}-\lambda_{1}\lambda+\sigma_{1}^{2}-\sigma_{2}^{2}=0$. With $\Delta:=\frac{(\sigma^{2}_{1}-\sigma^{2}_{2})}{\lambda_{1}^{2}}$, this gives
$\lambda_{M+N-1}(\mathbf{H})$ $=\frac{1}{2}\lambda_{1}\left[1-\sqrt{1-4\Delta}\right]=\frac{1}{2}\lambda_{1}\frac{4\Delta}{1+\sqrt{1-4\Delta}}.$ (46)
Hence, using the fact that $1\leq 1+\sqrt{1-4\Delta}\leq 2$ we obtain $\frac{(\sigma^{2}_{1}-\sigma^{2}_{2})}{\lambda_{1}}\leq\lambda_{M+N-1}(\mathbf{H})\leq\frac{2(\sigma^{2}_{1}-\sigma^{2}_{2})}{\lambda_{1}}$. To obtain the bounds for the condition numbers, by (46), we have
$\frac{(M+N)^{2}}{2M^{2}N^{2}(\sigma_{1}^{2}-\sigma_{2}^{2})}=\frac{\lambda_{1}^{2}}{2(\sigma^{2}_{1}-\sigma^{2}_{2})}\leq\kappa(\mathbf{H})\leq\frac{\lambda_{1}^{2}}{(\sigma^{2}_{1}-\sigma^{2}_{2})}=\frac{(M+N)^{2}}{M^{2}N^{2}(\sigma_{1}^{2}-\sigma_{2}^{2})}.$
Note the spectral gap of $\bm{\Pi}^{\top}\bm{\Pi}$ is $\rho(\bm{\Pi}^{\top}\bm{\Pi})=\sigma_{1}^{2}-\sigma_{2}^{2}$, then this gives the bounds in (45). ∎
@@
Let $M=N$ and $\bm{\Pi}$ be a positive coupling matrix with uniform marginals. The eigenvalues of $\mathbf{H}$ are
$\lambda_{j}(\mathbf{H})=\frac{1}{N}+\sigma_{j},\quad\lambda_{2N+1-j}(\mathbf{H})=\frac{1}{N}-\sigma_{j},\quad 1\leq j\leq N,$ (47)
where $\{\sigma_{j}\}$ are the singular values of $\bm{\Pi}$ in descending order. In particular, $\sigma_{1}=\frac{1}{N}$, $\lambda_{1}(\mathbf{H})=\frac{2}{N}$ and $\lambda_{2N}=0$. The condition number of $H(\bm{\Pi})$ is bounded by
$\frac{2}{N^{2}\rho(\bm{\Pi}^{\top}\bm{\Pi})}\leq\kappa(\mathbf{H})\leq\frac{4}{N^{2}\rho(\bm{\Pi}^{\top}\bm{\Pi})},$ (48)
where $\rho(\bm{\Pi}^{\top}\bm{\Pi})$ is spectral gap of the matrix $\bm{\Pi}^{\top}\bm{\Pi}$.
The Sinkhorn algorithm alternatively re-scales the rows and columns of the coupling matrix to achieve the marginal constraints. It produces a sequence of coupling matrices $\{\bm{\Pi}^{(l)}\}$ that converges to $\bm{\Pi}^{*}$ entry-wisely, i.e., $\lim_{l\rightarrow+\infty}\Pi_{ij}^{(l)}=\Pi_{ij}^{*}$ [31, 32, 23]. In practice, the Sinkhorn iteration stops when a criterion is met. One stopping criterion is that the marginal distributions of $\bm{\Pi}^{(l)}$ are entry-wise $\delta$ away from the given $\bm{\mu}$ and $\bm{\nu}$. Thus, an important question is whether the condition number of $\mathbf{H}$ of $\bm{\Pi}^{(l)}$ is controlled.
The next proposition shows that if $\delta=\|\bm{\Pi}-\bm{\Pi}^{*}\|_{F}$ is small with $\|\cdot\|_{F}$ denotes the Frobenius norm, the condition number of $\mathbf{H}$ of $\bm{\Pi}^{(l)}$ is almost as large as the condition number of $\mathbf{H}^{*}$. The proof is based on Weyl’s inequality, and we postpone it to supplementary materials.
Proposition 12 (Condition number of H matrices in Sinkhorn).
Let $\bm{\Pi}^{*}$, with uniform marginal distributions, be the optimal coupling matrix minimizing the EOT distance. Assume that the coupling matrix $\widehat{\bm{\Pi}}$ is computed by an early-stopped Sinkhorn algorithm that satisfies
@@
Then, the eigenvalues of $\mathbf{H}$ of $\widehat{\bm{\Pi}}$ satisfies
$|\lambda_{k}(\mathbf{H})-\lambda_{k}(\mathbf{H}^{*})|\leq\delta+\delta_{2},\quad 1\leq k\leq N+M.$ (50)
In particular, if $\delta+\delta_{2}=t\frac{MN}{M+N}\rho(\bm{\Pi}^{\top}\bm{\Pi})$ with $t\in[0,1)$, where $\sigma_{1},\sigma_{2}$ denoting the largest two singular values of $\bm{\Pi}^{*}$, the condition number of $\mathbf{H}$ of $\widehat{\bm{\Pi}}$ is bounded by
$\frac{1-t\Delta}{(2+t)\Delta}\leq\kappa(\mathbf{H})\leq\frac{1+t\Delta}{(1-t)\Delta},$
where $\Delta=(\frac{MN}{M+N})^{2}\rho(\bm{\Pi}^{\top}\bm{\Pi})$, while $\frac{1}{2\Delta}\leq\kappa(\mathbf{H}^{*})\leq\frac{1}{\Delta}$.
The above bounds apply to general $\mathbf{H}$ matrices in entropy-regularized Sinkhorn algorithms. However, these bounds do not show explicit dependence on $\epsilon$, the strength of regularization. In the next section, we study the dependence of the condition number on $\epsilon$ and $N$ for point-clouds datasets.
4.3 Ill-conditioned H matrices from data clouds
We investigate in this section the condition number of the $\mathbf{H}^{*}$ for a specific example of EOT that matches data clouds with $N$ points. In this simple setting, $M=N$ and both marginals are uniform, so the condition number of the $\mathbf{H}^{*}$ matrix is $\kappa=\frac{2}{N\lambda_{2N-1}}$ by Corollary 11. Therefore, it suffices to investigate the dependence of the smallest positive eigenvalue, $\lambda_{2N-1}$, on $N$ and $\epsilon$.
We show that the smallest positive eigenvalue of the $\mathbf{H}^{*}$ matrix can decay at rate $O(e^{-\frac{1}{\epsilon}})$ for a fixed $N$ and at $O(1/N)$ for a fixed $\epsilon$. These asymptotic results are proved for equally-spaced points on the unit circle in Example 13, and are numerically demonstrated for random data clouds sampled from a uniform distribution. So the $\mathbf{H}^{*}$ matrix in practice could be very ill-conditioned, and the proper regularization to resolve numerical instability is necessary.
@@ 4.3 Ill-conditioned H matrices from data clouds
Consider $N$ equally spaced points on the unit circle $\{\mathbf{y}_{i}=\left[\begin{array}[]{ccccccccc}\cos x_{i}&\sin x_{i}\end{array}\right]\}_{i=0}^{N-1}$, where $x_{i}=\frac{2\pi i}{N}$. Let $\bm{\mu}=\frac{1}{N}\mathbb{1}_{N}$ be the uniform distribution, we are interested in the spectrum of $\mathbf{H}^{*}$ associated to the symmetric entropic regularized optimal transport loss $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\mu})$. The coupling matrix is
$\bm{\Pi}^{*}=\underset{\bm{\Pi}\in\mathbf{U}(\bm{\mu},\bm{\mu})}{\operatorname{arg}\operatorname{min}}\;\sum_{i=1}^{N}\sum_{j=1}^{N}C_{ij}\Pi_{ij}+\epsilon\text{KL}(\bm{\Pi},{\bm{\mu}}\otimes\bm{\mu})$
with $\epsilon>0$, where $C_{ij}=\|\mathbf{y}_{i}-\mathbf{y}_{j}\|_{2}^{2}$. Denote its condition number by $\kappa(\mathbf{H}^{*})=\frac{\lambda_{1}(\mathbf{H}^{*})}{\lambda_{2N-1}(\mathbf{H}^{*})}$. Then, the following statements hold true.
@@
    $\bm{\Pi}^{*}=\frac{\mathbf{K}}{\lambda_{1}(\mathbf{K})N}$, where the Gibbs kernel $\mathbf{K}\in\mathbb{R}^{N\times N}$ is a symmetric matrix with entries $K_{ij}=\exp\left(-\frac{C_{ij}}{\epsilon}\right)$ and $\lambda_{1}(\mathbf{K})$ is the largest eigenvalue of $\mathbf{K}$.
@@
    The first two singular values of $\bm{\Pi}^{*}$ are $\sigma_{1}(\bm{\Pi}^{*})=\frac{1}{N}$, $\sigma_{2}(\bm{\Pi}^{*})=\frac{\lambda_{2}(\mathbf{K})}{\lambda_{1}(\mathbf{K})N}$, and the largest and smallest positive eigenvalues of $\mathbf{H}^{*}$ are $\lambda_{1}(\mathbf{H}^{*})=\frac{2}{N}$, and $\lambda_{2N-1}(\mathbf{H}^{*})=\frac{1}{N}-\sigma_{2}(\bm{\Pi}^{*})$, where $\lambda_{2}(\mathbf{K})$ is the second largest eigenvalue of $\mathbf{K}$.
@@
    The smallest positive eigenvalue of $\mathbf{H}^{*}$ and the condition number of $\mathbf{H}^{*}$ satisfies
    $\lim_{\epsilon\rightarrow 0^{+}}\lim_{N\rightarrow+\infty}\frac{N\cdot\lambda_{2N-1}(\mathbf{H}^{*})}{\epsilon}$ $=\frac{1}{4},\quad$ $\lim_{\epsilon\rightarrow 0^{+}}\lim_{N\rightarrow+\infty}\epsilon\cdot\kappa(\mathbf{H}^{*})$ $=8,$ (51)
    $\lim_{N\rightarrow+\infty}\lim_{\epsilon\rightarrow 0^{+}}\frac{N\lambda_{2N-1}(\mathbf{H}^{*})}{r_{N,\epsilon}}$ $=4\pi^{2},\quad$ $\lim_{N\rightarrow+\infty}\lim_{\epsilon\rightarrow 0^{+}}r_{N,\epsilon}\kappa(\mathbf{H}^{*})$ $=\frac{1}{2\pi^{2}},$ (52)
@@
We postpone the derivation to supplementary materials.
For a fixed $\epsilon$, when $N$ is large enough, the smallest positive eigenvalue $\lambda_{2N-1}$ scales as $\frac{\epsilon}{4N}$, which is numerically illustrated in Figure 1(a). Then the condition number scales as $\frac{8}{\epsilon}$. Meanwhile for fixed $N$, when $\epsilon$ is small enough (e.g., when $\epsilon<\frac{4\pi^{2}}{N^{2}}$), the smallest positive eigenvalue scales as $\frac{4\pi^{2}r_{N,\epsilon}}{N}$, which is numerically illustrated in Figure 1(b). Then the condition number scales as $\frac{1}{2\pi^{2}r_{N,\epsilon}}$, which grows exponentially at rate $O(e^{-\frac{1}{\epsilon}})$. For example, for $N=50$ and $\epsilon=0.0001$, the condition number of $\mathbf{H^{*}}$ is larger than $10^{70}$; in this case, a truncated SVD for $\mathbf{H}^{*}$ is crucial in calculating the Hessian of EOT and the gradient of Sinkhorn distance. On the other hand, when $N=1500$ and $\epsilon=0.0001$, the condition number of $\mathbf{H}^{*}$ is only about $8\times 10^{4}$. In addition, this scale phenomenon is also observed for some random datasets as well.
Figure 1: Decay of the smallest positive eigenvalue $\lambda_{2N-1}$ in $N$ and $\epsilon$. Equally spaced points on the unique circle: (a) $\lambda_{2N-1}\approx\frac{\epsilon}{4N}$ when $N>\frac{2\pi}{\sqrt{\epsilon}}$; (b) $\lambda_{2N-1}\approx 4\pi^{2}r_{N,\epsilon}$ when $\epsilon<\frac{4\pi^{2}}{N^{2}}$. Uniformly distributed point cloud in unit square $[0,1]^{2}$: (c) $\lambda_{2N-1}=O(\frac{1}{N})$ when $N$ is large; (d) $\lambda_{2N-1}=O(e^{-\frac{1}{\epsilon}})$ when $\epsilon$ is small.
Next, we further numerically investigate the case of point-clouds datasets that are sampled from the uniform distribution in the unit square $[0,1]^{2}$. Similarly, we are interested in the spectrum of $\mathbf{H}^{*}$ associated with the symmetric $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\mu})$. In Figure 1(d) show that $\lambda_{2N-1}=O(e^{-\frac{1}{\epsilon}})$ when $\epsilon$ is small enough for each fixed N, and Figure 1(c) show $\lambda_{2N-1}=O(\frac{1}{N})$ when $N$ is large for each fixed $\epsilon$. These asymptotic orders are the same as the analytical results proved in Example 13, but the exact limit depends on the distribution of data points, and it is beyond the scope of this study.
In summary, $\mathbf{H}^{*}$ can be severely ill-conditioned with the smallest eigenvalue at the order of $O(e^{-\frac{1}{\epsilon}})$ when $\epsilon$ is small, or $O(\frac{1}{N})$ when $N$ is large. Thus, when solving a linear system with $\mathbf{H}^{*}$, it is important to properly regularize the ill-posed inverse problem.